OpenStudy (anonymous):

what is the value of the discriminant of the equation x^2-9x-10=0

OpenStudy (anonymous):

The discriminant of \(ax^2+bx+c\) is \(\sqrt{b^2+4ac}\). Identify \(a,b,c\) for this quadratic and plug them in.

Directrix (directrix):

The discriminant is b² - 4ac.

Directrix (directrix):

The discriminant is b² - 4ac. b² - 4ac = (-9)² - 4*(1)*(-10) = ? @thanhuyenlebich

OpenStudy (anonymous):

@thanhuyenlebich, disregard the square root. My mistake. Thanks for clearing up @Directrix